# Illicit Promotion on Twitter

### v2

@@ 6.2. Cybercrime Operators
Scale of illicit promotion operators. In total , we have observed 474,979 distinct PIP campaigns. Among these campaign, 93.50% campaigns turn out to be singleton groups and each graph contains only an author node (i.e., a PIP account) and the tweets published by this account. These singleton groups constitute 73.9% PIPs and 76.5% PIP accounts. As for the the remaining 6.5% of the campaigns, 85% are campaigns with only one contact, 8.61% have two contacts, and 3.85% have three or more contacts. With respect to distribution across categories, 87.54% campaigns promote a single PIP category, 9.91% campaigns promote two categories and 2.55% campaigns promote three or more categories.
Representative 4 campaigns of illicit promotion. we looked further into the campaigns and discovered some commonly used schemes in cybercrime promotion. In general, the most notable characteristic of the campaigns is that most operators take control of at least two accounts to accomplish promotion and these accounts are connected by common contacts or websites. We present four campaigns as below under the name of Cluster-1 to Cluster-4.
Cluster-1. Figure7(a) visualizes cluster-1 as a graph wherein nodes denote either PIP accounts or contacts, and the size of each node represents the number of PIPs it is associated with. As shown in Figure7(a), cluster-1 consists of 3 nodes (two account nodes and one WeChat contact), 3 edges and 6,328 PIPs, and all PIPs are classified as data-leakage. The two Twitter accounts have published similar amounts of PIPs and promoted the same WeChat contact, and are very likely to be operated by the same person or organization.
(a) Cluster-1.
(b) Cluster-2.
(c) Cluster-3.
(d) Cluster-4.
Figure 7. Representative 4 PIP clusters.
7. Discussion
Recommendations for real-world PIP mitigation. We recommend OSNs should invest more efforts into detecting and removing PIPs from their platforms, particularly for PIPs not in English. Also, PIP detection should not limit to scrutinizing the text elements, but take into consideration all elements of an OSN post and its posting account. Also considering PIPs are operated as campaigns, clustering-based detection will be promising and efficient if a low fake alarm rate can be achieved. Also, as demonstrated in our evaluation, the series of tools developed in this study can benefit future endeavors to mitigate illicit promotion in online social networks. Besides, many PIPs prefer IM channels to interact with potential customers while promoting on OSNs, emphasizing the importance of cross-platform collaboration, especially between OSNs and IM platforms.
@@ 7. Discussion
Code and data release. To facilitate reproduction of our machine learning models (e.g., the binary PIP classifier and the NER-based contact extractor), the source code for training and testing will be released along with the respective ground truth datasets. Also, to avoid misuse by miscreants, the whole dataset of PIPs and thier contacts will be provided upon request and background checking.
8. Conclusion
Through this study, we have qualified and quantified the prevalence of diverse posts of illicit promotions (PIPs) on Twitter. Also, it is observed that illicit operators have adopted various evasion tactics when composing and distributing PIPs, which partially explains why so many PIPs could circumvent the content moderation of Twitter, get posted, and keep alive for months before being unpublished. What is also observed is that accounts of instant messaging platforms, especially end-to-end encrypted ones, are frequently used as the next hops for the communication between PIP victims and the underlying illicit operators. Such a cross-platform operation pattern also highlights the importance of security collaborations among OSN and instant messaging platforms for the mitigation of illicit promotion activities.
Appendix A Extra PIP Examples of Various Categories
Figure 8 presents typical examples of illicit promotion posts on Twitter.
@@ Appendix A Extra PIP Examples of Various Categories
Appendix B The List of Jargon Words as Observed from PIPs

### v3

@@ 6.2. Cybercrime Operators
Scale of illicit promotion operators. In total , we have observed 474,979 distinct PIP campaigns. Among these campaign, 93.50% campaigns turn out to be singleton groups and each graph contains only an author node (i.e., a PIP account) and the tweets published by this account. These singleton groups constitute 73.9% PIPs and 76.5% PIP accounts. As for the the remaining 6.5% of the campaigns, 85% are campaigns with only one contact, 8.61% have two contacts, and 3.85% have three or more contacts. With respect to distribution across categories, 87.54% campaigns promote a single PIP category, 9.91% campaigns promote two categories and 2.55% campaigns promote three or more categories. 4 representative clusters are shown in Appendix K.
7. Recommendations and Disclosure
Recommendations for real-world PIP mitigation. We recommend that OSNs should invest more efforts into detecting and removing PIPs from their platforms, particularly for PIPs not in English. PIP detection should not be limited to scrutinizing the text elements, but also take into consideration all elements of an OSN post and its posting account. Also considering PIPs are operated as campaigns, clustering-based detection will be promising and efficient if a low fake alarm rate can be achieved. Also, as demonstrated in our evaluation, the series of tools developed in this study can benefit future endeavors to mitigate illicit promotion in online social networks. Besides, many PIPs prefer IM channels to interact with potential customers while promoting on OSNs, emphasizing the importance of cross-platform collaboration, especially between OSNs and IM platforms.
@@ 7. Recommendations and Disclosure
Acknowledgements.
This work is jointly supported by the National Natural Science Foundation of China under Grant No. 62302473, and University of Science and Technology of China through the Innovation Fund for Young Investigators.
Appendix A Search poisoning attack
The search poisoning attack is a type of attack wherein the attacker compromises a legitimate website, injects promotional and harmful webpages, induces the search engines to index these webpages with high page ranks, and ultimately expose benign search users to the injected harmful webpages. Such promotional infections have been used to promote diverse cybercrime activities, e.g., online gambling (Liao et al., 2016a), unlicensed pharmacies (Leontiadis et al., 2014), etc. As revealed by (Leontiadis et al., 2014), the median time to recover such promotional infections is around 15 days. To detect promotional infections, John et al. (John et al., 2011) proposed the detection of infected webpages by looking at its URL parameters instead of visiting the webpage. Besides, Liao et al. (Liao et al., 2016a) utilized the semantic inconsistency between the injected cybercrime content and the legitimate context of the infected website to decide if a webpage is a promotional infection or not.
Appendix B The Binary PIP Classifier
B.1. Ground truth labeling process
An iterative process is followed for labeling, involving 1) searching Twitter with PIP-relevant keywords, which gives crawled tweets; 2) labeling a sampled subset of the crawled tweets to update ground-truth; 3) training a weak PIP classifier using the updated ephemeral ground-truth; 4) applying the weak PIP classifier to predict crawled tweets that are unlabeled, identifying both false positive and false negative predictions; and 5) updating ground-truth accordingly. Besides, when a sample is labeled, it is assigned with not only the binary PIP class, but also one of the PIP categories as listed in Table 2. This iterative process continues until no new PIP categories emerge, each PIP category is well represented in ground-truth and the PIP classifier has achieved a good performance when evaluated on the crawled tweets. Furthermore, inter-rater agreement rate over 90% is achieved across all labeling tasks. Particularly, for 1,000 samples from ground-truth, the agreement rate is 99.8%.
One thing to note, when composing non-PIP posts, we consider only PIP candidates that are not true PIPs, rather than using regular tweets. This is based on the observation that such PIP candidates tend to sit closer to the decision boundary than regular tweets and can help train a more robust PIP classifier. Besides, despite having more PIPs than non-PIPs in the ground truth, we don’t observe any negative impact on the binary classifier’s performance, which is well demonstrated by the evaluation of wild tweets and crawled tweets. Instead, we observe that fewer PIPs can yield comparable performance. For instance, by removing 4K PIPs for balance, a binary PIP classifier trained on this dataset achieved a precision of 96.6% and a recall of 97.4%. More PIPs, on the other hand, help enhance the ground-truth diversity in categories and languages, and are used to train/evaluate the multi-class PIP classifiers as introduced in §3.2.
B.2. Three-fold evaluation
Our three-fold evaluation consists of 5-fold cross-validation upon ground-truth, evaluation on crawled tweets, and evaluation on wild tweets. Table 7 lists the results of the five-fold cross validation. We can see that the text-only transformer-based classifier has achieved a performance that is comparable to that of the multi-modality one, and in the meantime better than that of the classic classifier.
Table 7. The performance of binary PIP classifiers.
Model Precision Recall F1-Score
Text-Only Transformer 97.25% 96.04% 96.64%
Multi-Modality Transformer 96.56% 97.43% 97.00%
SVM 95.29% 95.30% 95.29%
Random Forest 95.08% 94.77% 94.81%
Besides, the multi-modality model has the extra cost of downloading and preprocessing the involved media files, while classic algorithms are inferior to transformer models in terms of data efficiency, multilingual support, automatic feature engineering, and generalizability. We thus choose the text-only transformer model as the default binary PIP classifier.
We further evaluated the selected PIP classifier on unlabeled tweets as collected by our tweet crawler using PIP-relevant keywords. Given prediction results, 500 positive predictions were sampled out for manual validation along with 500 negative predictions. And the manual validation reveals a precision of 94.20% and a recall of 100%.
Then, to further evaluate the generalizability of our binary PIP classifier, we applied it to wild tweets, namely, multiple daily tweet snapshots from the Twitter Archiving Project [^3] https://archive.org/details/twitterarchive, which archives the tweet stream on a daily basis at a sampling ratio of 1%. Upon predictions on wild tweets, the same manual validation process was applied, which revealed a false positive rate (FPR) of 0.12% with a probability threshold of 0.5. The FRP can be further lowered by tuning the probability threshold. Furthermore, the precision was observed to be 96.8% while the recall is 99.4%. Therefore, considering the low FPR along with the high precision and recall, we can conclude with high confidence that this binary PIP classifier can be used independently to detect PIPs from wild tweets.
Appendix C Language Models
A language model (langmodel, 2023) is a probability distribution over a set of natural language words, e.g., RNN-based word2vec (Mikolov et al., 2013) and transformer-based BERT (Devlin et al., 2018). A language model is usually trained on a large corpus of unlabeled text documents through self-supervised training tasks such as masked language modeling, i.e., predicting a missing word in a given sentence. Since the hidden layers of such language models can well represent natural language words in a fixed-size and high-dimensional vector space and are thus commonly used for text encoding, i.e., word embedding. However, training a large and general-purpose language model is both time-consuming and computing-intensive, which motivates the emergence and increasing adoption of the paradigm of pre-training and fine-tuning. In this paradigm, to fulfill an NLP task, rather than building up a DNN model from scratch, a general-purpose and pre-trained language model is first adopted and then fine-tuned on a labeled dataset that is specific to the given NLP task and can be small-scaled. This paradigm outperforms numerous NLP tasks (Devlin et al., 2018; Koroteev, 2021; González-Carvajal and Garrido-Merchán, 2020). In this study, to facilitate the detection and understanding of PIPs, we have adopted the paradigm of pre-training and fine-tuning in multiple text classification tasks, e.g., binary PIP classification, multiclass PIP classification, and PIP contact recognition, as detailed in §3. Particularly, across these NLP tasks, we adopt the multilingual BERT (mul, 2023) as the pre-trained language model, which is built up through two self-supervised training tasks, namely, masked language modeling and next sentence prediction, upon unlabeled Wikipedia corpus which contains numerous text documents written in 102 natural languages.
Appendix D More Details on the Multi-Class PIP Classifier
D.1. The Basis for Defining PIP Categories
For each of the 10 categories, we have observed a reasonable volume of samples when labeling PIPs. We also define another category as others to denote PIPs that don’t fit in well for the aforementioned categories. Then, regarding the naming of these categories, we try our best to make them self-explained while keeping aligned with relevant terms used in previous works (Yang et al., 2017a; Yuan et al., 2018; Cao et al., 2012; Lin, 2011; Zhou et al., 2018; West and Bhattacharya, 2016). Particularly, 8 of the 10 categories are considered as illegal in Twitter safety and cybercrime rules (twi, 2023c). Two exceptions are gambling and surrogacy, likely because they vary significantly in legitimacy across jurisdictions. However, we decide to include them as PIP categories considering multiple factors. Particularly, promotion posts of both categories are often correlated with jurisdictions wherein they are illegal. For instance, 68.29% of surrogacy posts are in Chinese while surrogacy in China is prohibited. Besides, multiple previous studies on illicit promotion also consider both categories (Yang et al., 2017a; Yuan et al., 2018; Cao et al., 2012; Lin, 2011; Zhou et al., 2018; West and Bhattacharya, 2016).
D.2. More Evaluation Results
For the text-only classification, we still fine-tuned bert-base-multilingual-cased(mul, 2023), a multilingual transformer-based language model. For the multimodal classification, both the text and the visual modality for each PIP have been taken as the input. To build up such a model, ResNet-152 is used to embed the visual input (i.e., an image), and bert-base-multilingual-cased is used to embed the text input. For both models, 80% of the ground truth dataset is used for training, and the remaining 20% is held out for testing. As listed in Table 8, both have achieved very good performance across a set of well-acknowledged metrics. We choose the text-only model as the default multiclass PIP classifier.
Table 8. The performance of multiclass PIP classifiers.
Model Precision Recall F1-Score
Text-Only Transformer 98.82% 98.80% 98.80%
Multi-Modality Transformer 96.86% 97.73% 97.27%
Appendix E The PIP Contact Extractor
PIP contact extractor is a synthesized tool to extract both website URLs and IM accounts from PIP texts including QQ, Wechat, Telegram, Whatsapp and LINE. Among these contacts, website URLs can be easily identified through regular expression matching. However, many URLs deserve further processing for two factors. On the one hand, some URLs were found to redirect visitors to an IM account, e.g., a Telegram URL https://t.me/{account_id}, and we call such URLs as IM URLs. Such IM URLs will be further processed to extract the respective IM type and IM accounts, which is still achieved by defining and applying regular expressions specific to IM platforms. For instance, WhatsApp’s IM URL pattern is https://wa.me/{phone_number} while it is https://line.me/ti/p/{accountID} for LINE. On the other hand, due to the character limit for each tweet, many URLs were found to have been shortened through popular URL shorteners, e.g., https://bit.ly, and https://tinyurl.com. These shortened URLs will be further visited to recover the true URL of the final landing page. One thing to note, a LINE URL could also be shortened by http://lin.ee/XXXXXXX, and it is http://wa.link/XXXXX for Whatsapp. Then, after the true URL is recovered, the IM contacts will be extracted.
One thing to note, in some PIPs, contacts are also embedded in images or videos. However, meanwhile, we observe that most of these contacts are also promoted in the text profile of the respective PIP account. As our PIP hunter scans not only tweet contents, but also account profiles, such contacts can still be captured using our text-only classifiers and extractors. To extend coverage to corner cases where the contact is embedded only in images/videos, OCR can be first applied, and we leave it as future work to explore.
Appendix F PIP Categories
(a) A PIP with an image attached to signal child pornography.
(b) The account profile of the PIP presented in Figure 4(a).
Figure 4. An example of a pornography PIP.
(a) A drug promotion for methamphetamine in Chinese.
(b) A drug promotion for marijuana in Thai.
Figure 5. Examples of drug PIPs.
Among all categories, pornography is the most prevalent one, with 69.78% PIPs belonging to this category. To evade detection, many child pornography PIPs contain only innocent or seemingly harmless text. Instead, the images attached to these tweets contain necessary visual elements that both hint child pornography as well as providing explicit instructions regarding how to access child pornography. Additionally, geographical names are often used in PIPs so as to advertise location-based illegal services, e.g., local sex services. For example, the following two tweets from two different accounts, “Candice Bridges Elton CopperField #Guangzhou #Guangzhou Massage” and “Lesley Browne Baldwin Yerkes #Guangzhou #Guangzhou Massage”, both use massage and wellness as a way to promote illegal sex services in Guangzhou, China. Also, both differ only in the text but share the same images, hashtags, geographical names, and contacts.
Besides, 8.04% PIPs involve the promotion of various drugs, e.g., methamphetamine, marijuana, and heroin. As illustrated in Figure 5, these tweets promote drugs in Chinese and Thai, respectively. Also, these PIPs have utilized jargon words. For example, ”猪肉”, a Chinese word denoting pork, is used in illicit promotion to name methamphetamine. It is indeed surprising to see them get posted successfully on Twitter since even the tweet text itself appears sufficient to decide it is drug-related, not to mention the hashtags as well as the images. Besides, weapon PIPs often specify the detailed model of the weapons on sale. And some even include images to demonstrate the item’s availability. Similarly, harassment PIPs are mainly used to promote harassment as a service, such as SMS bombing, and phone call bombing. Lastly, PIPs in the category of data theft&leakage encompass many products, such as hacking services, stolen account credentials, unauthorized access to confidential information, and personal data leakage.
Figure 6 shows more concrete PIP examples.
@@ Appendix F PIP Categories
Appendix G The List of Jargon Words as Observed from PIPs
@@
Appendix H Evasion Techniques
Below, we present the different types of evasion techniques that are commonly adopted by miscreants to avoid Twitter content moderation.
Jargon words. Across PIPs of different categories, jargon words are commonly used. Particularly, through looking into PIPs, we have manually identified a set of 108 different jargon words that are embedded into 31.21% of all PIPs, which can only serve as a lower-bound estimate for the adoption of jargon words in PIPs. Also, the adoption of jargon words is observed for all the illicit categories. For example, in Chinese PIPs of drugs, ”叶子” (leaves in English), or its emoji is used to refer to marijuana, while ”猪肉” (pork) and ”冰”(ice) and their emojis are used to represent methamphetamine. These terms are derived from either the color or the shape of the respective illegal items. Additionally, a metaphor of the nature of the activity is also popular as jargon words for PIPs. For instance, the farmer and its emoji are used in drug PIPs to refer to people who plant marijuana. The use of jargon words helps illicit promotion operators blend their content with benign tweets, thus impeding content moderation. More jargon examples can be found in Appendix G.
Appendix I Social engagement on PIPs
Over time, PIPs accumulate engagements such as likes, replies, retweets and quotes (i.e. retweets with comments). Given the date a PIP is published $t_{p}$ and the date it is crawled $t_{c}$, $t_{e}=t_{c}-t_{p}$ is the elapse time of a PIP, i.e., the days passed since it is posted. As described in §5, 90% PIPs survive the first two months, thus we group PIPs based on elapse time ranging from 1 day to 60 days, and calculate the average engagement of each group. As shown in Figure 7, PIPs indeed receive a non-negligible volume of engagement, e.g., PIPs can receive 374 likes on average after being posted for 30 days, while the average number of likes a regular tweet can receive is 37 (ave, 2021). What is also observed is that the extent of engagement declines gradually after 35 days. We believe it is because PIPs with more engagements are more likely to be detected and unpublished by Twitter, which renders the survivorship bias where PIPs with fewer engagements are more likely to survive longer and be counted when calculating engagement of long elapse time. However we are not clear about the reason why average retweets show significant fluctuations after 40 days.
Figure 7. Average engagements of PIPs over time.
Appendix J Additional analysis on PIP Contacts
J.1. Evasion techniques for promoting PIP contacts
As discussed in §5, PIP operators have adopted various evasion or promotion tactics so as to make the resulting PIPs appear benign. Furthermore, various evasion techniques have also been observed to hide a contact deeply in the respective PIP. Equipped with such evasion tactics, it would become challenging for the OSN platform to extract the contacts from a detected PIP. For instance, some miscreants use code words, emojis, abbreviations, or unconventional formatting to obscure contact information. Also, contact information of some PIPs are hidden in images or even videos.
Visual representation and steganography. The visual representation of contacts involves presenting contact information as visual elements in images or even videos. This can be further strengthened by the use of human-written text (Figure 5(a)), which makes it difficult for even OCR systems to recognize and extract the respective contact information. Embedding contact information within seemingly innocuous elements is also a way of stealthy promotion, e.g., embedding the contact in the poll options instead of the tweet’s main text. Additionally, illicit operators may also embed contact information within any frames of innocuous videos. Currently, our PIP contact extractor supports only the retrieval of contacts from text inputs, namely, the PIP text and the profile text of a PIP account. We leave it as our future work to extract contacts from media files and other seemingly innocuous components of a PIP.
Private messaging. By requesting potential buyers to send private messages for more information (e.g., Figure 6(a) in Appendix LABEL:appendix:cp_examples), illicit operators can avoid the direct exposure of their contact information in OSNs. This tactic enables them to selectively respond to potential buyers and increase the effectiveness of their promotional efforts.
J.2. Contact preference across PIP categories
Regarding contact selection, illicit promotion campaigns of different categories exhibit distinct preferences. Table 10 presents, for each PIP category, how the extracted contact entities distribute across well-known contact types. Porn, money laundering, and weapon sales predominantly prefer Telegram accounts, which are associated with 43.60%, 68.17%, 73.36% of PIPs of the corresponding category, respectively. For harassment activities, QQ emerges as the most favored contact method, accounting for 56.23% of harassment PIPs. Besides, for data theft & leakage and fake documents, WeChat is much more popular than others.
Category1 Telegram Wechat QQ LINE WA2
Pornography 43.6% 30.24% 34.02% 2.49% 0.02%
Illegal Drug 36.99% 24.91% 13.79% 14.73% 1.12%
Gambling 42.45% 41.39% 24.73% 2.96% 0.97%
Surrogacy 22.92% 67.76% 14.26% 0.03% 1.38%
Harassment 25.60% 22.78% 56.23% 0.01% 0.00%
Money Laundering 68.17% 19.77% 17.91% 0.04% 0.24%
Weapon Sales 73.36% 18.80% 7.86% 0.1% 2.5%
DTL 2 35.75% 58.91% 11.51% 0.02% 1.50%
FFD 2 16.64% 68.76% 14.81% 0.47% 0.00%
Crowdturfing 55.23% 44.18% 11.04% 0.00% 0.00%
  • 1
    Note that one PIP may include more than one contact.
  • 2
    WA denotes WhatsApp, FFD is short for Forgery and Fake Documents and DTL is short for Data Theft and Leakage.
J.3. Threat alerts from VirusTotal
To profile the threats of websites embedded in PIPs, we retrieved and analyzed their threat reports from VirusTotal (vir, 2023), a widely recognized open threat exchange platform. Due to the rate limit of VT, we only randomly sampled 20k PIP website URLs (corresponding to 7,391 FQDNs) for analysis. As shown in Table 11, 0.80% URLs have one or more alarms triggered, while it is 2.46% for FQDNs. Furthermore, among these alarmed URLs, 11.38% are considered malware websites, while 24.39% are considered phishing websites.
Table 11. Statistics of VirusTotal reports for PIP websites.
Category Count Reported Alarmed Malware Phishing
URLs 20,000 7.88% 0.80% 0.07% 0.15%
FQDNs 7,391 69.30% 2.46 % 0.32% 0.70%
Appendix K Representative Campaigns Underpinning PIPs
we looked further into the campaigns and discovered some commonly used schemes in cybercrime promotion. In general, the most notable characteristic of the campaigns is that most operators take control of at least two accounts to accomplish promotion and these accounts are connected by common contacts or websites. We present four campaigns as below under the name of Cluster-1 to Cluster-4.
Cluster-1. Figure8(a) visualizes cluster-1 as a graph wherein nodes denote either PIP accounts or contacts, and the size of each node represents the number of PIPs it is associated with. As shown in Figure8(a), cluster-1 consists of 3 nodes (two account nodes and one WeChat contact), 3 edges, and 6,328 PIPs, and all PIPs are classified as data-leakage. The two Twitter accounts have published similar amounts of PIPs and promoted the same WeChat contact, and are very likely to be operated by the same person or organization.
(a) Cluster-1.
(b) Cluster-2.
(c) Cluster-3.
(d) Cluster-4.
Figure 8. Representative 4 PIP clusters.